cream has approximately 22% of butterfat.how many gallons of cream must be mixed with 2% milk to give 20 gallons of milk containing 4% of butterfat
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

cream has approximately 22% of butterfat.how many gallons of cream must be mixed with 2% milk to give 20 gallons of milk containing 4% of butterfat
-------------------------
Equation:
fat + fat = fat
0.22x + 0.02(20-x) = 0.04*20
-----
Multiply thru by 100 to get:
22x + 2*20 - 2x = 4*20
20x = 2*20
x = 2 gallons (amt. of cream needed in the mixture)
=============================
Cheers,
Stan H.

Answer by mananth(16946)

cream has approximately 22% of butterfat.how many gallons of cream must be mixed with 2% milk to give 20 gallons of milk containing 4% of butterfat
cream 22% fat quantity x
milk fat 2% quantity (20-x)
Blend 4% quantity = 20
concentration of cream fat + concentration of milk fat = concentration of butter fat
0.22x +0.02(20-x) = 20*0.04
0.22x + 0.4 -0.02x=0.8
0.20x = 0.8 -0.4
0.2x= 0.4
x= 2 gallons of cream